A line of slope 1/2 intersects a line of slope 3. Find the size of the acute angle that these lines form, to the nearest tenth of a degree.
Answer by stanbon(75887)

A line of slope 1/2 intersects a line of slope 3. Find the size of the acute angle that these lines form, to the nearest tenth of a degree.
---------------------------------------------------------
Make of sketch of y = (1/2)x and of y = 3x
-------------------------------------------------
y = (1/2)x has points (x,x/2)
tan(angle made with positive x axis) = (1/2)/1 = 1/2
Angle made with positive x axis = tan^-1(1/2) = 26.5 degrees
-----------------------------------------------------------------
y = 3x has points (x,3x)
t(angle made with positive x axis) = 3
Angle made with positive x axis = tan^-1(3) = 71.57
-------------------------------------------------------------------
Angle between the lines is 71.57-26.5 = 45.1 degrees
===================================================================
Cheers,
Stan H.